OpenStudy (anonymous):

Evaluate: sin(Arcsec 1)

OpenStudy (anonymous):

Arcsec 1 = 0 sin 0 = 0

OpenStudy (anonymous):

arcsec x = 1/cos x

OpenStudy (anonymous):

sec x = 1/cos x

myininaya (myininaya):

hamed is right! :)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

damn, i said sec so arcsec is actually meant to be ???

OpenStudy (nikita2):

sec(arcsec1)= 1 => 1/cos(arcsec1) = 1 => cos(arcsec1)=1 => sin(arcsec1) = 0 (for example because sin^2 + cos^2 = 1)

myininaya (myininaya):

heres the other 2 traingles i figure more examples the same way would help